OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (-1, 4) and is parallel to the line y = 2x - 3?

OpenStudy (anonymous):

Parallel lines have the same gradient so the gradient of the line y= 2x- 3 is 2. Now use this formula 'y- y1= m(x- x1)' where y1= 4, m= 2 and x1= -1 :) hope u understand!

OpenStudy (anonymous):

so it would be 2x + y = -6 or 2x - y = -6 @KissMyAxe

OpenStudy (anonymous):

It's wrong

OpenStudy (anonymous):

*Your calculation